OpenStudy (anonymous):

Will award medal: For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 2 x - y = -6 A) -6 - x B) -6 + x C) 6 - x D) 6 + x

OpenStudy (anonymous):

2x + y = 2 x - y = -6

OpenStudy (texaschic101):

x - y = -6 -y = -x - 6 -- multiply by -1 to make y positive y = x + 6 or y = 6 + x

OpenStudy (texaschic101):

answer D

OpenStudy (anonymous):

Thank you. ^^

OpenStudy (texaschic101):

anytime :)